OpenStudy (anonymous):

Line 1 (strawberries) : (1,2) , (2,4) Line 2 (blueberries) : (1,1.5) , (2,3) Write an equation in slope intercept form to represent the cost of the blueberries. @ttb123456789 @swagmaster47 Can you help?

OpenStudy (anonymous):

@ttb123456789

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

WELL LE ME THINK

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=ax+b. ok so now do the following. (1.5-3)/(1-2)=1.5 is the slope a=1.5 so now y=1.5x+b. Now substitute for example (2, 3) and count: 3=1.5*2+b, b=3-3=0. Can you do the following?

OpenStudy (anonymous):

i got the slope of 1.5, is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

so wtite now the qation for me i want to check

OpenStudy (anonymous):

y=1.5x+5?

OpenStudy (anonymous):

uh no

OpenStudy (anonymous):

1.5* not 5

OpenStudy (anonymous):

y=1.5x+b

OpenStudy (anonymous):

did you saw the value of b in my solving?

OpenStudy (anonymous):

the y-interseptio if wrong it should be 0

OpenStudy (anonymous):

y=1.5x+0?

OpenStudy (anonymous):

yes but it is wwritten as just y=1.5x

OpenStudy (anonymous):

it is when the y-interseption is 0

OpenStudy (anonymous):

Yay. Thank you! You're awesome!